# Quantum Mycielskians: symmetries, twin vertices and distinguishing labelings

### v1

@@ 2. Quantum isomorphisms of Mycielskians
$\rho_{\mathcal{F},\mathcal{G}}(\xi)=p(\xi\otimes 1).$
We define a map $\widehat{p}\in\operatorname{B}(L^{2}(\mu_{r-1}(\mathcal{G})),L^{2}(\mu_{r-1}(\mathcal{F})))\otimes\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))$:
$\widehat{p}=\varpi_{\mathcal{F},0}\varpi_{\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\mathcal{F},k}p\varpi_{\mathcal{G},k}^{*},$
where $\varpi_{\mathcal{G},j}=\iota_{\mathcal{G},j}\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))}$ and $\varpi_{\mathcal{F},j}=\iota_{\mathcal{F},j}\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))}$, with $\iota_{\mathcal{G},j}$ and $\iota_{\mathcal{F},j}$ denoting the isometric embeddings of $L^{2}(\mathcal{G})$ and $L^{2}(\mathcal{F})$ into $L^{2}(\mu_{r-1}(\mathcal{G}))$ and $L^{2}(\mu_{r-1}(\mathcal{F}))$, respectively.
@@
$\widehat{p}\widehat{p}^{*}=\varpi_{\mathcal{F},0}\varpi_{\mathcal{F},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\mathcal{F},k}pp^{*}\varpi_{\mathcal{F},k}^{*}=\sum\limits_{j=0}^{r}\varpi_{\mathcal{F},j}\varpi_{\mathcal{F},j}^{*}=\mathrm{id}_{L^{2}(\mu_{r-1}(\mathcal{F}))},$
$\widehat{p}^{*}\widehat{p}=\varpi_{\mathcal{G},0}\varpi_{\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{k,\mathcal{G}}pp^{*}\varpi_{\mathcal{G},k}^{*}=\sum\limits_{j=0}^{r}\varpi_{\mathcal{G},j}\varpi_{\mathcal{G},j}^{*}=\mathrm{id}_{L^{2}(\mu_{r-1}(\mathcal{G}))},$
@@
$\widehat{p}A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}$ $=\biggl{(}\varpi_{\mathcal{F},0}\varpi_{\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\mathcal{F},k}p\varpi_{\mathcal{G},k}^{*}\biggr{)}\biggl{(}\delta\varpi_{\mathcal{G},r}\eta_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},0}^{*}+\delta\varpi_{\mathcal{G},0}\eta^{*}_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},r}^{*}+\varpi_{\mathcal{G},1}A_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},1}^{*}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum\limits_{k=1}^{r-1}\bigl{(}\varpi_{\mathcal{G},k}A_{\mathcal{G}}\varpi_{\mathcal{G},k+1}^{*}+\varpi_{\mathcal{G},k+1}A_{\mathcal{G}}\varpi_{\mathcal{G},k}^{*}\bigr{)}\biggr{)}$
$=\delta\varpi_{\mathcal{F},0}\eta_{\scriptscriptstyle\mathcal{G}}^{*}\varpi_{\mathcal{G},r}^{*}+\delta\varpi_{\mathcal{F},r}p\eta_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},0}^{*}+\varpi_{\mathcal{F},1}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},1}^{*}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum\limits_{j=1}^{r-1}\bigl{(}\varpi_{j,\mathcal{F}}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},j+1}^{*}+\varpi_{\mathcal{F},j+1}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\mathcal{G},j}^{*}\bigr{)},$
@@
$A_{\mu_{r-1}(\mathcal{F})}\widehat{p}=\delta\varpi_{\mathcal{F},0}\eta_{\mathcal{F}}^{*}p\varpi_{\mathcal{G},r}^{*}+\delta\varpi_{\mathcal{F},r}\eta_{\mathcal{F}}\varpi_{\mathcal{G},0}^{*}$ $+\varpi_{\mathcal{F},1}A_{\mathcal{F}}p\varpi_{\mathcal{G},1}^{*}$
$+\sum\limits_{j=1}^{r-1}\bigl{(}\varpi_{j,\mathcal{F}}A_{\mathcal{F}}p\varpi_{\mathcal{G},j+1}^{*}+\varpi_{\mathcal{F},j+1}A_{\mathcal{F}}p\varpi_{\mathcal{G},j}^{*}\bigr{)}.$
Since $p\eta_{\scriptscriptstyle\mathcal{G}}=\eta_{\mathcal{F}}$, we also have $\eta_{\scriptscriptstyle\mathcal{G}}^{*}=\eta_{\mathcal{F}}^{*}p$. This shows that indeed the intertwining property holds. Finally, the map
$\rho_{\mu_{r-1}(\mathcal{F}),\mu_{r-1}(\mathcal{G})}\left(\left[\begin{smallmatrix}\lambda\\ \xi_{1}\\[-3.0pt] \vdots\\[1.0pt] \xi_{r}\end{smallmatrix}\right]\right)=\widehat{p}\left(\left[\begin{smallmatrix}\lambda\\ \xi_{1}\\[-3.0pt] \vdots\\[1.0pt] \xi_{r}\end{smallmatrix}\right]\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{G}),\operatorname{QAut}(\mathcal{F}))}\right)$
is a unital *-homomorphism by construction. ∎
@@ 3. Quantum symmetries of $\mu_{r-1}(\mathcal{G})$ from quantum symmetries of $\mathcal{G}$
Let $\mathcal{G}$ be a quantum graph. For $r\geq{2}$ define the linear map $\Phi_{r}\colon\operatorname{C}(\mu_{r-1}(\mathcal{G}))\to\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ by
@@
To prove the multiplicativity of $\Phi_{r}$ we temporarily denote by $m^{(0)}$, $m^{(1)}$, $m^{(2)}$ and $m^{(3)}$ the multiplication maps on $\mathbb{C}$, $\operatorname{C}(\mathcal{G})$, $\operatorname{C}(\mu_{r-1}(\mathcal{G}))$ and $\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ respectively (the last one defined only on the algebraic tensor product) and recall from [5, Section 4] that
@@
Hence for $(p,q)\neq(0,0)$, with $s=\max\{p,q\}$ and $\boldsymbol{\chi}$ denoting the flip map, by (2) we have
[…]
@@
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\eta_{\mathcal{G}}\iota_{0}^{*}+\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}\eta_{\mathcal{G}}^{*}\iota_{r}^{*}+\iota_{1}A_{\scriptscriptstyle\mathcal{G}}\iota_{1}^{*}+\sum_{k=1}^{r-1}(\iota_{k}A_{\scriptscriptstyle\mathcal{G}}\iota_{k+1}^{*}+\iota_{k+1}A_{\scriptscriptstyle\mathcal{G}}\iota_{k}^{*}),$
where $\eta_{\mathcal{G}}$ is the unit map $\mathbb{C}\to\operatorname{C}(\mathcal{G})$. It follows from this equation that
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{0}$ $=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\eta_{\mathcal{G}},$
@@
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{r}$ $=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}\eta_{\mathcal{G}}^{*}+\iota_{r-1}A_{\scriptscriptstyle\mathcal{G}}.$
@@
$(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})$ $\Phi_{r}\bigl{(}\iota_{0}(\lambda)\bigr{)}=A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{0}(\lambda)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\bigl{(}\eta_{\mathcal{G}}(\lambda)\bigr{)}\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$
@@
for $l=2,\dotsc,r-1$, and finally using the fact that $\eta_{\mathcal{G}}^{*}=\psi$ and that $\psi$ is $\rho_{\scriptscriptstyle\mathcal{G}}$-invariant we arrive at
@@
$=\delta_{\scriptscriptstyle\mathcal{G}}(\iota_{0}\eta_{\mathcal{G}}^{*}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)+(\iota_{r-1}A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
@@
$=\delta_{\scriptscriptstyle\mathcal{G}}\bigl{(}\iota_{0}(\eta_{\mathcal{G}}^{*}a)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\bigr{)}+(\iota_{r-1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)$
$=\Phi_{r}\bigl{(}\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}(\eta_{\mathcal{G}}^{*}a)+\iota_{r-1}(a)\bigr{)}=\Phi_{r}\bigl{(}A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{r}(a)\bigr{)}.$
@@
We can amplify actions of quantum groups on $\operatorname{C}(\mathcal{G})$ into those on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ for a Hilbert space $\mathscr{H}$. In fact, the action $\rho_{\scriptscriptstyle\mathcal{G}}$ of $\operatorname{QAut}(\mathcal{G})$ on $\mathcal{G}$ defines
$\rho_{\mathcal{G},\mathscr{H}}=(\mathrm{id}_{\operatorname{C}(\mathcal{G})}\otimes\boldsymbol{\chi})(\rho_{\mathcal{G}}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})})\colon\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\longrightarrow\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G})),$
the action of $\operatorname{QAut}(\mathcal{G})$ on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$. For a partition of unity $\mathfrak{P}\subset\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$, we consider the action $\bigl{(}\rho_{\mathcal{G},\mathscr{H}}\bigr{)}_{\mathfrak{P}}$ of $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$.

### v2

@@ 2. Quantum isomorphisms of Mycielskians
$\rho_{\scriptscriptstyle\mathcal{F},\mathcal{G}}(\xi)=p(\xi\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))}).$
We define an element $\widehat{p}\in\operatorname{B}(L^{2}(\mu_{r-1}(\mathcal{G})),L^{2}(\mu_{r-1}(\mathcal{F})))\otimes\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))$:
$\widehat{p}=\varpi_{\scriptscriptstyle\mathcal{F},0}\varpi_{\scriptscriptstyle\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\scriptscriptstyle\mathcal{F},k}p\varpi_{\scriptscriptstyle\mathcal{G},k}^{*},$
where $\varpi_{\scriptscriptstyle\mathcal{G},j}=\iota_{\scriptscriptstyle\mathcal{G},j}\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))}$ and $\varpi_{\scriptscriptstyle\mathcal{F},j}=\iota_{\scriptscriptstyle\mathcal{F},j}\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))}$, with $\iota_{\scriptscriptstyle\mathcal{G},j}$ and $\iota_{\scriptscriptstyle\mathcal{F},j}$ ($j=1,\ldots,r$)denoting the isometric embeddings of $L^{2}(\mathcal{G})$ and $L^{2}(\mathcal{F})$ into $L^{2}(\mu_{r-1}(\mathcal{G}))$ and $L^{2}(\mu_{r-1}(\mathcal{F}))$, respectively.
@@
$\widehat{p}\widehat{p}^{*}$ $=\varpi_{\scriptscriptstyle\mathcal{F},0}\varpi_{\scriptscriptstyle\mathcal{F},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\scriptscriptstyle\mathcal{F},k}pp^{*}\varpi_{\scriptscriptstyle\mathcal{F},k}^{*}=\sum\limits_{j=0}^{r}\varpi_{\scriptscriptstyle\mathcal{F},j}\varpi_{\scriptscriptstyle\mathcal{F},j}^{*}=\mathrm{id}_{L^{2}(\mu_{r-1}(\mathcal{F}))},$
$\widehat{p}^{*}\widehat{p}$ $=\varpi_{\scriptscriptstyle\mathcal{G},0}\varpi_{\scriptscriptstyle\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\scriptscriptstyle\mathcal{G},k}pp^{*}\varpi_{\scriptscriptstyle\mathcal{G},k}^{*}=\sum\limits_{j=0}^{r}\varpi_{\scriptscriptstyle\mathcal{G},j}\varpi_{\scriptscriptstyle\mathcal{G},j}^{*}=\mathrm{id}_{L^{2}(\mu_{r-1}(\mathcal{G}))},$
@@
$\widehat{p}A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}$ $=\biggl(\varpi_{\scriptscriptstyle\mathcal{F},0}\varpi_{\scriptscriptstyle\mathcal{G},0}^{*}+\sum\limits_{k=1}^{r}\varpi_{\scriptscriptstyle\mathcal{F},k}p\varpi_{\scriptscriptstyle\mathcal{G},k}^{*}\biggr)\biggl(\delta\varpi_{\scriptscriptstyle\mathcal{G},r}\eta_{\scriptscriptstyle\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},0}^{*}+\delta\varpi_{\scriptscriptstyle\mathcal{G},0}\eta^{*}_{\scriptscriptstyle\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},r}^{*}+\varpi_{\scriptscriptstyle\mathcal{G},1}A_{\scriptscriptstyle\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},1}^{*}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum\limits_{k=1}^{r-1}\bigl(\varpi_{\scriptscriptstyle\mathcal{G},k}A_{\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},k+1}^{*}+\varpi_{\scriptscriptstyle\mathcal{G},k+1}A_{\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},k}^{*}\bigr)\biggr)$
$=\delta\varpi_{\scriptscriptstyle\mathcal{F},0}\eta_{\scriptscriptstyle\mathcal{G}}^{*}\varpi_{\scriptscriptstyle\mathcal{G},r}^{*}+\delta\varpi_{\scriptscriptstyle\mathcal{F},r}p\eta_{\scriptscriptstyle\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},0}^{*}+\varpi_{\scriptscriptstyle\mathcal{F},1}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},1}^{*}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\sum\limits_{j=1}^{r-1}\bigl(\varpi_{\scriptscriptstyle\mathcal{F},j}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},j+1}^{*}+\varpi_{\scriptscriptstyle\mathcal{F},j+1}pA_{\scriptscriptstyle\mathcal{G}}\varpi_{\scriptscriptstyle\mathcal{G},j}^{*}\bigr),$
@@
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{F})}\widehat{p}=\delta\varpi_{\scriptscriptstyle\mathcal{F},0}\eta_{\scriptscriptstyle\mathcal{F}}^{*}p\varpi_{\scriptscriptstyle\mathcal{G},r}^{*}+\delta\varpi_{\scriptscriptstyle\mathcal{F},r}\eta_{\scriptscriptstyle\mathcal{F}}\varpi_{\scriptscriptstyle\mathcal{G},0}^{*}$ $+\varpi_{\scriptscriptstyle\mathcal{F},1}A_{\scriptscriptstyle\mathcal{F}}p\varpi_{\scriptscriptstyle\mathcal{G},1}^{*}$
$+\sum\limits_{j=1}^{r-1}\bigl(\varpi_{\scriptscriptstyle\mathcal{F},j}A_{\scriptscriptstyle\mathcal{F}}p\varpi_{\scriptscriptstyle\mathcal{G},j+1}^{*}+\varpi_{\scriptscriptstyle\mathcal{F},j+1}A_{\scriptscriptstyle\mathcal{F}}p\varpi_{\scriptscriptstyle\mathcal{G},j}^{*}\bigr).$
Since $p\eta_{\scriptscriptstyle\mathcal{G}}=\eta_{\scriptscriptstyle\mathcal{F}}$, we also have $\eta_{\scriptscriptstyle\mathcal{G}}^{*}=\eta_{\scriptscriptstyle\mathcal{F}}^{*}p$. This shows that indeed the intertwining property holds. Finally, the map
$\rho_{\scriptscriptstyle\mu_{r-1}(\mathcal{F}),\mu_{r-1}(\mathcal{G})}\left(\left[\begin{smallmatrix}\lambda\\ \xi_{1}\\[-3.0pt] \vdots\\[1.0pt] \xi_{r}\end{smallmatrix}\right]\right)=\widehat{p}\left(\left[\begin{smallmatrix}\lambda\\ \xi_{1}\\[-3.0pt] \vdots\\[1.0pt] \xi_{r}\end{smallmatrix}\right]\otimes\mathds{1}_{\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))}\right)$
is a unital *-homomorphism $\operatorname{C}(\mu_{r-1}(\mathcal{G}))\to\operatorname{C}(\mu_{r-1}\mathcal{F}))\otimes\mathcal{O}(\operatorname{QAut}(\mathcal{F}),\operatorname{QAut}(\mathcal{G}))$ by construction. Since the algebra $\mathcal{O}(\operatorname{QAut}(\mu_{r-1}(\mathcal{F})),\operatorname{QAut}(\mu_{r-1}(\mathcal{G})))$ is universal for such homomorphisms, it has a non-zero homomorphic image, so it is itself non-zero. It follows that $\mu_{r-1}(\mathcal{G})$ and $\mu_{r-1}(\mathcal{F})$ are quantum isomorphic. ∎
@@ 3. Quantum symmetries of $\mu_{r-1}(\mathcal{G})$ from quantum symmetries of $\mathcal{G}$
Let $\mathcal{G}$ be a quantum graph. Then $\operatorname{QAut}(\mathcal{G})$ acts faithfully on its quantum Mycielskians. More precisely, for $r\geq{2}$ define the linear map $\Phi_{r}\colon\operatorname{C}(\mu_{r-1}(\mathcal{G}))\to\operatorname{C}(\mu_{r-1}(\mathcal{G}))\otimes\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ by
@@
To prove the multiplicativity of $\Phi_{r}$ we temporarily denote by $m^{(0)}$, $m^{(1)}$, $m^{(2)}$ and $m^{(3)}$ the multiplication maps on $\mathbb{C}$, $\operatorname{C}(\mathcal{G})$, $\operatorname{C}(\mu_{r-1}(\mathcal{G}))$ and $\operatorname{C}(\operatorname{QAut}(\mathcal{G}))$ respectively (the last one defined only on the algebraic tensor product) and recall from [6, Section 4] that
@@
Hence for $(p,q)\neq(0,0)$, with $\boldsymbol{\chi}$ denoting the flip map, by (2) we have
$\begin{aligned} \Phi_{r}&\circ m^{(2)}\bigl(\iota_{p}(a)\otimes\iota_{q}(b)\bigr)=\Phi_{r}\left(\delta_{p,q}\sqrt{\tfrac{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}{\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}\iota_{p}(ab)\right)\\ &=\delta_{p,q}\sqrt{\tfrac{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}{\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}(\iota_{p}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(ab)\\ &=\delta_{p,q}\sqrt{\tfrac{1+r\delta_{\scriptscriptstyle\mathcal{G}}^{2}}{\delta_{\scriptscriptstyle\mathcal{G}}^{2}}}(\iota_{p}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(m^{(1)}\otimes m^{(3)})(\mathrm{id}_{\operatorname{C}(\mathcal{G})}\otimes\boldsymbol{\chi}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\rho_{\scriptscriptstyle\mathcal{G}}\otimes\rho_{\scriptscriptstyle\mathcal{G}})(a\otimes b)\\ &=\delta_{p,q}(m^{(2)}\otimes m^{(3)})(\iota_{p}\otimes\iota_{p}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\mathrm{id}_{\operatorname{C}(\mathcal{G})}\otimes\boldsymbol{\chi}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\rho_{\scriptscriptstyle\mathcal{G}}\otimes\rho_{\scriptscriptstyle\mathcal{G}})(a\otimes b)\\ &=\delta_{p,q}(m^{(2)}\otimes m^{(3)})(\mathrm{id}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))}\otimes\boldsymbol{\chi}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\iota_{p}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\otimes\iota_{p}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\rho_{\scriptscriptstyle\mathcal{G}}\otimes\rho_{\scriptscriptstyle\mathcal{G}})(a\otimes b)\\ &=(m^{(2)}\otimes m^{(3)})(\mathrm{id}_{\operatorname{C}(\mu_{r-1}(\mathcal{G}))}\otimes\boldsymbol{\chi}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})(\Phi_{r}\otimes\Phi_{r})\bigl(\iota_{p}(a)\otimes\iota_{q}(b)\bigr)\end{aligned}$
@@
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\eta_{\scriptscriptstyle\mathcal{G}}\iota_{0}^{*}+\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}\eta_{\scriptscriptstyle\mathcal{G}}^{*}\iota_{r}^{*}+\iota_{1}A_{\scriptscriptstyle\mathcal{G}}\iota_{1}^{*}+\sum_{k=1}^{r-1}(\iota_{k}A_{\scriptscriptstyle\mathcal{G}}\iota_{k+1}^{*}+\iota_{k+1}A_{\scriptscriptstyle\mathcal{G}}\iota_{k}^{*}),$
where $\eta_{\scriptscriptstyle\mathcal{G}}$ is the unit map $\mathbb{C}\to\operatorname{C}(\mathcal{G})$. It follows from this equation that
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{0}$ $=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\eta_{\scriptscriptstyle\mathcal{G}},$
@@
$A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{r}$ $=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}\eta_{\scriptscriptstyle\mathcal{G}}^{*}+\iota_{r-1}A_{\scriptscriptstyle\mathcal{G}}.$
@@
$(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})$ $\Phi_{r}\bigl(\iota_{0}(\lambda)\bigr)=A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{0}(\lambda)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}=\delta_{\scriptscriptstyle\mathcal{G}}\iota_{r}\bigl(\eta_{\scriptscriptstyle\mathcal{G}}(\lambda)\bigr)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}$
@@
for $l=2,\dotsc,r-1$, and finally using the fact that $\eta_{\scriptscriptstyle\mathcal{G}}^{*}=\psi$ and that $\psi$ is $\rho_{\scriptscriptstyle\mathcal{G}}$-invariant we arrive at
@@
$=\delta_{\scriptscriptstyle\mathcal{G}}(\iota_{0}\eta_{\scriptscriptstyle\mathcal{G}}^{*}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)+(\iota_{r-1}A_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(a)$
@@
$=\delta_{\scriptscriptstyle\mathcal{G}}\bigl(\iota_{0}(\eta_{\scriptscriptstyle\mathcal{G}}^{*}a)\otimes\mathds{1}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))}\bigr)+(\iota_{r-1}\otimes\mathrm{id}_{\operatorname{C}(\operatorname{QAut}(\mathcal{G}))})\rho_{\scriptscriptstyle\mathcal{G}}(A_{\scriptscriptstyle\mathcal{G}}a)$
$=\Phi_{r}\bigl(\delta_{\scriptscriptstyle\mathcal{G}}\iota_{0}(\eta_{\scriptscriptstyle\mathcal{G}}^{*}a)+\iota_{r-1}(a)\bigr)=\Phi_{r}\bigl(A_{\scriptscriptstyle\mu_{r-1}(\mathcal{G})}\iota_{r}(a)\bigr).$
@@
We can amplify actions of quantum groups on $\operatorname{C}(\mathcal{G})$ into those on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ for a Hilbert space $\mathscr{H}$. In fact, the action $\rho_{\scriptscriptstyle\mathcal{G}}$ of $\operatorname{QAut}(\mathcal{G})$ on $\mathcal{G}$ defines an action $\rho_{\scriptscriptstyle\mathcal{G},\mathscr{H}}$ of $\operatorname{QAut}(\mathcal{G})$ on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$ by acting on $\operatorname{B}(\mathscr{H})$ trivially, that is $\rho_{\scriptscriptstyle\mathcal{G},\mathscr{H}}=(\mathrm{id}_{\operatorname{C}(\mathcal{G})}\otimes\boldsymbol{\chi})(\rho_{\scriptscriptstyle\mathcal{G}}\otimes\mathrm{id}_{\operatorname{B}(\mathscr{H})})$, where $\boldsymbol{\chi}$ is the flip map. For a partition of unity $\mathfrak{P}\subset\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$, we consider the action $\bigl(\rho_{\scriptscriptstyle\mathcal{G},\mathscr{H}}\bigr)_{\mathfrak{P}}$ of $\operatorname{QAut}(\mathcal{G})_{\mathfrak{P}}$ on $\operatorname{C}(\mathcal{G})\otimes\operatorname{B}(\mathscr{H})$.